OpenStudy (anonymous):

A wheel with a diameter of 3 feet makes a revolution every 2 minutes. How many feet will the wheel travel in 30 minutes? (A) 3π (B) 6π (C) 45π (D) 30π (E) 15π

jimthompson5910 (jim_thompson5910):

The diameter is 3 ft. So the radius is half that at 3/2 = 1.5 ft. The circumference (the distance around the wheel) is then C = 2pi*r = 2*pi*1.5 = 3pi feet. So one revolution (or complete 360 degree turn) will mean that the wheel has traveled 3pi feet. This one revolution occurs every 2 minutes. Since there are 30/2 = 15 two minute periods in 30 minutes, this means that the wheel has gone 15*3pi = 45 pi feet in 30 minutes. So the answer is C)

OpenStudy (anonymous):

Circumference: \[3\pi\] \[15\times 3\pi=45\pi\]

OpenStudy (anonymous):

thx